OpenStudy (anonymous):

If a(x) and b(x) are linear functions with one variable, which of the following expressions produces a quadratic function? a. (ab)(x) b. (a/b)(x) c. (a – b)(x) d. (a + b)(x)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle a(x)=mx+b }\) \(\large\color{black}{ \displaystyle b(x)=cx+a }\) (m, b, c, & a are just some constants) you need to produce a quadratic function: Quadratic function is in the form of: \(\large\color{black}{ \displaystyle y=ax^2+bx+c}\)

OpenStudy (solomonzelman):

So, which of the following operations do you think would make a second degree (quadratic) function